Adam borrowed $1,750 from Eve on November 28, 2019 and agreed to repay the debt with simple interest at the rate of 3.4% p.a. on July 15, 2020.

a. How much interest will Adam owe when the loan matures on July 15, 2020? __________

b. What is he maturity value (future value) of the loan on July 15, 2020? _________

Deadline 2 hours


1
Expert's answer
2021-02-18T13:07:55-0500


Where I is interest P is principal amount R is rate and T is period of time

P=1750

R=3.4

T = 8 months = 8/12= 2/3


"Interest= 1750×\frac{2}{3}×\frac{3.4}{100}=38.53"


"Future" "Value =1750+38.53= 1788.53"